# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ 3.2. OOD Detection Approaches with Both ID and OOD Data
Overview. In some known deployment scenarios, real OOD data can be easily collected at a low cost. Some methods based on this assumption focus on how to use OOD data for better detection performance. Differing from methods involving OOD Synthesis, in these research directions, models have access to real-world OOD data during the training phase. The primary focus in such problems is on optimizing the model’s decision boundary, rather than the OOD data itself. Due to the introduction of real OOD information, the boundary of these two categories will be accurately calculated.
Boundary Regularization. The Boundary Regularization class of methods belongs to the traditional Outlier Exposure (OE) approaches. The central idea of  Hendrycks et al. (2019) and  Hein et al. (2019) are to fully leverage OOD data to optimize the model’s decision boundary, thus achieving OOD detection. Proponents of this concept can utilize auxiliary anomaly datasets to enhance the OOD detector, enabling it to generalize and detect anomalous information not encountered during training. The central idea of this method can be grasped from Fig.  6.
Figure 6. Illustration of boundary regularization approaches. The key to model optimization strategy lies in the decision boundary of the model classification. Boundary regularization utilizes existing known ID data (the colored dots) and some OOD data (black dots) to optimize the potential decision boundary, compacting its classification boundary with abnormal data as tightly as possible without affecting the classification of ID data.
@@
over the parameters of $f$, where $x^{\prime}$ represents auxiliary anomaly data and $U$ denotes the uniform distribution. $\mathcal{L}_{\text{OE}}$ represents the cross-entropy loss with respect to $U$. The fundamental purpose is to compel the model to optimize the OOD data distribution to a uniform distribution, a principle that is universal in OE-type approaches. The specific design of $\mathcal{L}_{\text{OE}}$ can depend on other task requirements and the chosen OOD score. This design can utilize a maximum softmax probability baseline  (Hendrycks and Gimpel, 2017) detector to detect anomalous data. Compared to traditional softmax scores, EnergyOE  (Liu et al., 2020) builds upon OE by leveraging energy scores for better discrimination between ID and OOD samples, and it is less prone to issues of overconfidence. Specifically, its calculation formula:
@@
Outlier Mining. The traditional OE concept assumes the existence of ID input $\mathcal{D}_{\text{in}}$ and OOD input $\mathcal{D}_{\text{out}}$, both independently and heterogeneously distributed, originating from different sources. However, this premise cannot be fully guaranteed in the current training process due to potential noise in the training OOD data. Outlier Mining differs slightly from the traditional OE approach in that, although it also utilizes real-world OOD samples to address the issue, it focuses on identifying the optimal selection within the existing OOD data. The main process is depicted in Fig.  7.
Figure 7. Illustration of outlier mining approaches. Due to the limitation of whether real OOD data exists, methods like anomaly mining allow the model to access real OOD data during training. The black dots in the figure represent real OOD data sampled from other datasets, which do not overlap with the ID categories. Utilizing such OOD data to extrapolate more representative OOD data (orange dots) is a solution to this type of problem.
@@
Imbalance. Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information. The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two. To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024) incorporates a learnable tail class prototype during the training process. This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.  Choi et al. (2023) take a different approach, suggesting that factors affecting the performance of OOD detection may be related to the imbalance in the cross-class distribution of auxiliary OOD data. Consequently, it proposes an energy regularization loss, specifically regularizing auxiliary samples from the majority classes to address the issue of class imbalance in OOD data. EAT  (Wei et al., 2024) utilizes dynamically assigned virtual labels during the model training process to train OOD data, thereby expanding the classification space.
@@ 4. Problem: Training-agnostic OOD Detection
Training-agnostic OOD detection places a primary emphasis on adaptation strategies during test time, as opposed to the focus on classifier performance seen in training-driven OOD detection. Based on whether they rely on dependencies among test data, methods are categorized into two types: post-hoc and test-time adaptive approaches, as illustrated in Fig. 8. Post-hoc methods compute results for individual samples independently, unaffected by changes in other samples. In contrast, test-time adaptive methods exhibit dependencies among test samples.
Figure 8. Illustration of training-agnostic OOD detection approaches. Both methods require access to a pre-trained model. Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing. “A’/B’ ” means that the original features are deformed; Bank means that some samples are stored; Clustering means that clustering is performed for ID images.
@@ 4.1. Post-hoc OOD Detection Approaches
Post-hoc approaches are categorized into five types: Output-based, Distance-based, Gradient-based, Feature-based and Density-based. Recent work on this type of problem has some recent progress. A summary of the key factors involved in such methods is given in Table  1.
@@
Method Type Space
@@
MSP (Hendrycks and Gimpel, 2017) Output-Based $\checkmark$
@@
Mahalanobis (Lee et al., 2018b) Distance-Based $\checkmark$
@@
Grad (Lee and AlRegib, 2020) Gradient-Based $\checkmark$
@@
ODIN (Liang et al., 2018a) Feature-Based $\checkmark$ $\checkmark$ $\checkmark$
@@
BLOOD (Jelenić et al., 2024) $\checkmark$
SCALE (Xu et al., 2024) $\checkmark$
ConjNorm (Peng et al., 2024) Density-based $\checkmark$ $\checkmark$
@@
Output-based. Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others. MSP  (Hendrycks and Gimpel, 2017) was the first to employ the maximum softmax value to validate OOD detection effectiveness. For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category. In contrast to the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by comparing the maximum logit value in the logits vector output by the neural network. Logits, representing the model’s confidence in each category, are the neural network’s output before the softmax layer, without undergoing softmax transformation. Meanwhile, Energy  (Liu et al., 2020) introduces the Helmholtz Free Energy, which theoretically aligns with the input probability density and is less susceptible to overconfidence issues. GEN  (Liu et al., 2023b) introduces the concept of generalized entropy and directly utilizes Bregman divergence to compute the statistical distance between the model’s probability output and uniform distribution, aiming to identify OOD data. Additionally, leveraging sufficient prior knowledge might be a viable solution. ZODE  (Xue et al., 2024) perform predictions on samples across multiple pre-trained models simultaneously to determine whether multiple models can identify OOD samples, using this as a basis to distinguish between data. LogicOOD  (Kirchheim et al., 2024) present a novel approach that uses first-order logic for knowledge representation to perform OOD detection. This reasoning system uses prior knowledge to infer whether an input is consistent with the prior knowledge about the training distribution. It is particularly user-friendly in terms of interpretability, as it allows for comparing the output of samples against a knowledge base to determine whether they belong to the ID data.
@@
Distance-based. Another approach in OOD detection research focuses on measuring statistical distance metrics. Mahalanobis  (Lee et al., 2018b) is typically computed by calculating the distance between the feature vector and its mean. Specifically, for each class, we compute the mean and covariance matrix of its feature vectors. During testing, it calculates the Mahalanobis distance between the feature vector and the mean of each class. SSD  (Sehwag et al., 2021) essentially utilizes the Mahalanobis distance. After being trained on unlabeled ID data via self-supervised representation learning, it employs the Mahalanobis distance as a statistical measure for classification using the pre-trained model. In comparison, while the Mahalanobis distance makes strong distributional assumptions about the data, KNN  (Sun et al., 2022) explores the effectiveness of non-parametric nearest-neighbor distance for OOD detection. By measuring the k-nearest neighbor distance between input embeddings and training set embeddings, a threshold is designed to determine whether the data belongs to the ID data. NNGuide  (Park et al., 2023a) takes a step further in the direction of granularity by combining the idea of KNN. It assigns weights before the traditional OOD Score, depending on the nearest neighbor distance between the sample and the embeddings in the training set.
@@
Gradient-based. Grad  (Lee and AlRegib, 2020) research indicates that gradient-based methods also contribute to OOD detection by quantifying model uncertainty through the gradients propagated in reverse during backpropagation. If the input sample is an ID sample, the gradients of the model with respect to these samples tend to be relatively small and stable. Conversely, the gradients are typically larger or more irregular for OOD samples. GradNorm  (Huang et al., 2021) posits that the gradient magnitude for ID data surpasses that of OOD data. Leveraging this observation, it employs gradient vector norms, computing KL divergence through softmax output and uniform distribution backpropagation to detect OOD data. In contrast, GradOrth  (Behpour et al., 2023) adopts an alternative perspective, recognizing that crucial features of OOD data reside in a low-rank subspace. Consequently, it shifts focus to calculating gradient projection norms in this subspace to identify OOD data. GAIA  (Chen et al., 2023) employs a combination of gradient anomaly checks and aggregation. It enables the model to interpret indeterminacy attributively, introducing channel-wise average abnormality and zero-deflation abnormality without prior knowledge to gauge the extent of data distribution variations. The OPNP  (Chen et al., 2024a) method discovers that the OOD detection capability of the model is highly sensitive to parameters and activated neurons that are close to zero. Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability.
@@
where the parameter $\varepsilon$ is the perturbation magnitude. For a given input $x$, compute its logit output $p_{c}(x)$. Therefore, ReAct  (Sun et al., 2021) focuses on the high activation values in the intermediate results of the model. These activation values do not affect the model’s classification, but truncating high activation values can significantly improve OOD detection performance. VRA  (Xu et al., 2023), an extended iteration of ReAct  (Sun et al., 2021), builds upon the premise of ReAct which truncates only high activation values. However, VRA posits that this might not be the optimal solution and, therefore, employs a variational approach to seek an optimal solution. It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data. SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison. It employs a storage mechanism to calculate the average logits for each class, preparing this simple storage pattern for subsequent OOD detection tasks. Drawing inspiration from experience, DDCS  (Yuan et al., 2024) adapts to select suitable channels for data classification after correcting each channel in the neural network. These channels are evaluated based on inter-class similarity and variance to measure their discriminative power for ID data. LINe  (Yong Hyun Ahn, 2023) similarly emphasizes neuron outputs at the feature level. It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs.
@@
Feature shaping is emerging as a prominent area of research focus. This methodology entails the refinement of intermediate representations—specifically, the intermediate features—within the forward propagation phase of the model. The salient advantage of this approach lies in its non-disruptive nature concerning the original classification results, coupled with its simplicity and efficacy. Building upon this foundation, ViM  (Wang et al., 2022c) simultaneously considers the roles of features, logits, and probabilities, constructing virtual logits to aggregate information for decision-making. By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection. Its computation formula can be expressed in the following form:
(5) $-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
@@
where $\alpha$ is a scaling constant, computed by the model. Here $z=z^{P}+z^{P^{\perp}}$ and $z^{P^{\perp}}$ is the projection of $z$ to $P^{\perp}$. And it have $Wz^{P^{\perp}}=0$. Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020), and $f(z)$ represents the logit output of the model. The first term here represents virtual logits, while the second term represents the score of the energy function. Neco  (Ammar et al., 2024) subsequently reveals the prevalent phenomenon of neural collapse in contemporary neural networks, impacting OOD detection performance. The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data. ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted. NAC  (Liu et al., 2024a) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data. By quantifying this statistical property, NAC aims to distinguish between ID and OOD data. Based on the existing literature on feature shaping,  Zhao et al. (2024) utilize a piecewise constant shaping function to partition the feature domain into disjoint intervals, estimating a scalar within each interval as an approximation. As the interval width approaches zero, an approximation of the maximum logits can be obtained. The goal of BLOOD  (Jelenić et al., 2024) is to smooth the representations of intermediate layers in neural networks for predicting OOD data. This study found that, compared to OOD data, ID data exhibits smoother variations in the intermediate layer representations of neural networks. Leveraging this characteristic, new statistical measures can be designed to discriminate against anomalous data. In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH. SCALE  (Xu et al., 2024) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data. Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.

### v3

@@ 3.2. OOD Detection Approaches with Both ID and OOD Data
Overview. In some known deployment scenarios, real OOD data can be easily collected at a low cost. Some methods based on this assumption focus on how to use OOD data for better detection performance. Differing from methods involving OOD Synthesis, in these research directions, models have access to real-world OOD data during the training phase. The primary focus in such problems is on optimizing the model’s decision boundary, rather than the OOD data itself. Due to the introduction of real OOD information, the decision boundary between ID and OOD data is further accurately calculated.
Boundary Regularization. The Boundary Regularization class of methods belongs to the traditional Outlier Exposure (OE) approaches. The central idea of  Hendrycks et al. (2019) and  Hein et al. (2019) are to fully leverage OOD data to optimize the model’s decision boundary, thus achieving OOD detection. Proponents of this concept can utilize auxiliary anomaly datasets to enhance the OOD detector, enabling it to generalize and detect anomalous information not encountered during training. The central idea of this method can be grasped from Fig.  3 (e).
@@
over the parameters of $f$, where $x^{\prime}$ represents auxiliary anomaly data and $U$ denotes the uniform distribution. $\mathcal{L}_{\text{OE}}$ represents the cross-entropy loss with respect to $U$. Here, $\mathcal{D}_{\text{in}}$ denotes the distribution of ID data, while $\mathcal{D}_{\text{out}}^{\text{OE}}$ denotes the distribution of OOD.
The fundamental purpose is to compel the model to optimize the OOD data distribution to a uniform distribution, a principle that is universal in OE-type approaches. The specific design of $\mathcal{L}_{\text{OE}}$ can depend on other task requirements and the chosen OOD score. This design can utilize a maximum softmax probability baseline  (Hendrycks and Gimpel, 2017) detector to detect anomalous data. Compared to traditional softmax scores, EnergyOE  (Liu et al., 2020a) builds upon OE by leveraging energy scores for better discrimination between ID and OOD samples, and it is less prone to issues of overconfidence. Specifically, its calculation formula:
@@
Outlier Mining. The traditional OE concept assumes the existence of ID input $\mathcal{D}_{\text{in}}$ and OOD input $\mathcal{D}_{\text{out}}$, both independently and heterogeneously distributed, originating from different sources. However, this premise cannot be fully guaranteed in the current training process due to potential noise in the training OOD data. Outlier Mining differs slightly from the traditional OE approach in that, although it also utilizes real-world OOD samples to address the issue, it focuses on identifying the optimal selection within the existing OOD data. The main process is depicted in Fig.  3 (d).
@@
Long-tail ID data. Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information. The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two. To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024a) incorporates a learnable tail class prototype during the training process. This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.  Choi et al. (2023) take a different approach, suggesting that factors affecting the performance of OOD detection may be related to the imbalance in the cross-class distribution of auxiliary OOD data. Consequently, it proposes an energy regularization loss, specifically regularizing auxiliary samples from the majority classes to address the issue of class imbalance in OOD data. EAT  (Wei et al., 2024) utilizes dynamically assigned virtual labels during the model training process to train OOD data, thereby expanding the classification space.
@@ 4. Problem: Training-agnostic OOD Detection
Training-agnostic OOD detection focuses on adaptation strategies at test time for models that already possess downstream classification capabilities, as opposed to the focus on classifier performance seen in training-driven OOD detection. Based on whether they rely on dependencies among test data, methods are categorized into two types: post-hoc and test-time adaptive approaches, as illustrated in Fig. 4. Post-hoc methods compute results for individual samples independently, unaffected by changes in other samples. In contrast, test-time adaptive methods leverage information shared among test samples to enhance OOD detection.
Figure 4. Illustration of training-agnostic OOD detection approaches. Both methods require access to a pre-trained model. Post-hoc approaches do not involve any operations during the post-training phase, while test-time adaptive approaches necessitate adaptation based on the samples encountered during testing. “A’/B’ ” means that the original features are deformed; “Bank” means that some samples are stored; “Clustering” means that clustering is performed for ID images.
@@ 4.1. Post-hoc OOD Detection Approaches
Post-hoc approaches are categorized into five types: Output-based, Distance-based, Gradient-based, Feature-based and Density-based. Recent work on this type of problem has some recent progress. A summary of the key factors involved in such methods is given in Table  2.
@@
Type Method Space
@@
Output-Based MSP (Hendrycks and Gimpel, 2017) $\checkmark$
@@
Distance-Based Mahalanobis (Lee et al., 2018b) $\checkmark$
@@
Gradient-Based Grad (Lee and AlRegib, 2020) $\checkmark$
@@
Feature-Based ODIN (Liang et al., 2018a) $\checkmark$ $\checkmark$ $\checkmark$
@@
SCALE (Xu et al., 2024a) $\checkmark$
Density-Based ConjNorm (Peng et al., 2024) $\checkmark$ $\checkmark$
@@
Output-based. Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others. MSP  (Hendrycks and Gimpel, 2017) is the first to employ the maximum softmax value to validate OOD detection effectiveness. For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category. Departing from the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by utilizing the maximum value of the logits vector generated by the neural network. These logits constitute the raw, pre-Softmax layer outputs of the neural network, representing the model’s confidence in each category without undergoing Softmax transformation. Alternatively, the Energy  (Liu et al., 2020a) method defines an energy score conceptually derived from Helmholtz Free Energy. This score can theoretically reflect the input probability density. By assigning lower energy scores to ID data, this method distinguishes them from anomalous data. Additionally, it demonstrates reduced susceptibility to model overconfidence issues.
GEN  (Liu et al., 2023b) introduces the concept of generalized entropy and directly utilizes Bregman divergence to compute the statistical distance between the model’s probability output and uniform distribution, aiming to identify OOD data. In addition, leveraging sufficient prior knowledge might be a viable solution. ZODE  (Xue et al., 2024) performs predictions on samples across multiple pre-trained models simultaneously to determine whether multiple models can identify OOD samples, using this as a basis to distinguish between data. LogicOOD  (Kirchheim et al., 2024) presents a novel approach that uses first-order logic for knowledge representation to perform OOD detection. This reasoning system uses prior knowledge to infer whether an input is consistent with prior knowledge about the training distribution. It is particularly user-friendly in terms of interpretability, as it allows comparing the output of samples against a knowledge base to determine whether they belong to the ID data.
@@
Distance-based. Another approach in OOD detection research focuses on measuring statistical distance metrics. Mahalanobis  (Lee et al., 2018b) is typically computed by calculating the distance between the feature vector and its mean. Specifically, for each class, we compute the mean and covariance matrix of its feature vectors. During testing, it calculates the Mahalanobis distance between the feature vector and the mean of each class. Mahalanobis++ (Mueller and Hein, 2025b) also uses Mahalanobis distance, but applies L2 norm normalization to the features in the early stages of calculation to eliminate the instability in feature extraction across different models, which affects the distribution assumption they rely on. SSD  (Sehwag et al., 2021) essentially utilizes the Mahalanobis distance. After being trained on unlabeled ID data via self-supervised representation learning, it employs the Mahalanobis distance as a statistical measure for classification using the pre-trained model. In comparison, while the Mahalanobis distance makes strong distributional assumptions about the data, KNN  (Sun et al., 2022) explores the effectiveness of non-parametric nearest-neighbor distance for OOD detection. By measuring the k-nearest neighbor distance between input embeddings and training set embeddings, a threshold is designed to determine whether the data belongs to the ID data. NNGuide  (Park et al., 2023a) takes a step further in the direction of granularity by combining the idea of KNN. It assigns weights before the traditional OOD Score, depending on the nearest neighbor distance between the sample and the embeddings in the training set.
@@
Gradient-based. Grad  (Lee and AlRegib, 2020) research indicates that gradient-based methods also contribute to OOD detection by quantifying model uncertainty through the gradients propagated in reverse during backpropagation. If the input sample is an ID sample, the gradients of the model with respect to these samples tend to be relatively small and stable. Conversely, the gradients are typically larger or more irregular for OOD samples. GradNorm  (Huang et al., 2021) posits that the gradient magnitude for ID data surpasses that of OOD data. Leveraging this observation, it employs gradient vector norms, computing KL divergence through softmax output and uniform distribution backpropagation to detect OOD data. In contrast, GradOrth  (Behpour et al., 2023) adopts an alternative perspective, recognizing that crucial features of OOD data reside in a low-rank subspace. Consequently, it shifts focus to calculating gradient projection norms in this subspace to identify OOD data. GAIA  (Chen et al., 2023) employs a combination of gradient anomaly checks and aggregation. It enables the model to interpret indeterminacy attributively, introducing channel-wise average abnormality and zero-deflation abnormality without prior knowledge to gauge the extent of data distribution variations. The OPNP  (Chen et al., 2024a) method discovers that the OOD detection capability of the model is highly sensitive to parameters and activated neurons that are close to zero. Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability. In recent research, PRO (Chen et al., 2025a) introduces an adversarial score, with the core idea being to perturb the input data using gradient descent to search for the local minimum score near the original input. This method effectively lowers the confidence of OOD samples and is compatible with existing MSP detection methods during score calculation, thereby enhancing detection performance. S&I (Zhang et al., 2025a) identifies the issue of insufficient sample gradient feature discrimination in this type of method. By introducing adversarial samples and integrating gradients along the attribution path, it achieves a more accurate feature distinction strategy.
@@
where the parameter $\varepsilon$ is the perturbation magnitude. For a given input $x$, compute its logit output $p_{c}(x)$. Therefore, ReAct  (Sun et al., 2021) focuses on the high activation values in the intermediate results of the model. These activation values do not affect the model’s classification, but truncating high activation values can significantly improve OOD detection performance. VRA  (Xu et al., 2023), an extended iteration of ReAct  (Sun et al., 2021), builds upon the premise of ReAct which truncates only high activation values. However, VRA posits that this might not be the optimal solution and, therefore, employs a variational approach to seek an optimal solution. It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data. KANs (Canevaro et al., 2025) also focus on the neuron activation states. Due to the local neuroplasticity characteristic of Kolmogorov-Arnold Networks (Liu et al., 2025b), KANs distinguish between ID and OOD data by comparing the activation response differences between the trained and untrained KANs. SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison. It employs a storage mechanism to calculate the average logits for each class, preparing this simple storage pattern for subsequent OOD detection tasks. Drawing inspiration from experience, DDCS  (Yuan et al., 2024) adapts to select suitable channels for data classification after correcting each channel in the neural network. These channels are evaluated based on inter-class similarity and variance to measure their discriminative power for ID data. LINe  (Yong Hyun Ahn, 2023) similarly emphasizes neuron outputs at the feature level. It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs. CADRef (Ling et al., 2025) decouples the features by using a strategy based on the sign alignment between the relative features and model weights, dividing the sample features into positive and negative errors. Intuitively, the positive and negative error components influence the final output logits, which helps in identifying OOD images from this perspective. ITP (Xu and Yang, 2025) optimizes the model by quantifying the contribution of model parameters to ID data prediction, specifically by evaluating the contribution of each parameter through its partial derivatives and removing those with low contribution. Additionally, during the testing phase, ITP uses a right-tail Z-score test to assess whether any parameter exhibits overconfidence in the classification of a given test sample, further refining the model’s decision-making.
@@
Feature shaping is emerging as a prominent area of research focus. This methodology entails the refinement of intermediate representations—specifically, the intermediate features—within the forward propagation phase of the model. The salient advantage of this approach lies in its non-disruptive nature concerning the original classification results, coupled with its simplicity and efficacy. Building upon this foundation, ViM  (Wang et al., 2022c) simultaneously considers the roles of features, logits, and probabilities, constructing virtual logits to aggregate information for decision-making. By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection. Its computation of this score $S$ can be expressed in the following form:
(5) $S=-\alpha\|z^{P^{\perp}}\|^{2}+\text{LogSumExp}f(z),$
@@
where $\alpha$ is a scaling constant, computed by the model. Here $z=z^{P}+z^{P^{\perp}}$ and $z^{P^{\perp}}$ is the projection of $z$ to $P^{\perp}$. And it have $Wz^{P^{\perp}}=0$. Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020a), and $f(z)$ represents the logit output of the model. The first term here represents virtual logits, while the second term represents the score of the energy function. Neco  (Ammar et al., 2024) subsequently reveals the prevalent phenomenon of neural collapse in contemporary neural networks, impacting OOD detection performance. The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data. Similarly, NCI (Liu and Qin, 2023) studies the phenomenon of neural collapse and discovers that during training, ID samples tend to cluster around class weight vectors in the penultimate feature space, while OOD samples do not exhibit this clustering behavior. Therefore, NCI leverages the proximity between features and class weight vectors to effectively distinguish ID samples from OOD samples, while maintaining low computational cost. ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted. NAC (Liu et al., 2024b) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data. By quantifying this statistical property, NAC aims to distinguish between ID and OOD data. Based on the existing literature on feature shaping,  Zhao et al. (2024) utilize a piecewise constant shaping function to partition the feature domain into disjoint intervals, estimating a scalar within each interval as an approximation. As the interval width approaches zero, an approximation of the maximum logits can be obtained. The goal of BLOOD  (Jelenić et al., 2024) is to smooth the representations of intermediate layers in neural networks for predicting OOD data. This study found that, compared to OOD data, ID data exhibits smoother variations in the intermediate layer representations of neural networks. Leveraging this characteristic, new statistical measures can be designed to discriminate against anomalous data. In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH. SCALE  (Xu et al., 2024a) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data. Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.